OpenStudy (anonymous):

What is the equation of the line passing through (1, 7) and parallel to the line y = –x + 5 in standard form?

OpenStudy (anonymous):

plz someone help me

OpenStudy (anonymous):

two parallel lines have same slope. therefore, the slope of the required equation is -1.

OpenStudy (anonymous):

then use point slope form of an equation of a line.

OpenStudy (anonymous):

y-7=-1(x-1) y-7=-x+1 y=-x+8 x+y-8=0

OpenStudy (anonymous):

x+y-8=0 is the standard form of the equation